# LoopSR: Looping Sim-and-Real for Lifelong Policy Adaptation of Legged Robots

### v2

@@ Abstract
Reinforcement Learning (RL) has shown its remarkable and generalizable capability in legged locomotion through sim-to-real transfer. However, while adaptive methods like domain randomization are expected to enhance policy robustness across diverse environments, they potentially compromise the policy’s performance in any specific environment, leading to suboptimal real-world deployment due to the No Free Lunch theorem. To address this, we propose LoopSR, a lifelong policy adaptation framework that continuously refines RL policies in the post-deployment stage. LoopSR employs a transformer-based encoder to map real-world trajectories into a latent space and reconstruct a digital twin of the real world for further improvement. Autoencoder architecture and contrastive learning methods are adopted to enhance feature extraction of real-world dynamics. Simulation parameters for continual training are derived by combining predicted values from the decoder with retrieved parameters from a pre-collected simulation trajectory dataset. By leveraging simulated continual training, LoopSR achieves superior data efficiency compared with strong baselines, yielding eminent performance with limited data in both sim-to-sim and sim-to-real experiments.
@@ I Introduction
Reinforcement Learning (RL) has seen significant progress as a viable alternative for robotic control, but applying it to real-world scenarios remains challenging. Advances in simulation systems [1, 2] and extensive parallel training [3] have enhanced policy performance and robustness, but the sim-to-real gap persists.
Adaptive methods such as domain randomization [4, 5, 6, 7, 8] are widely used to address the gap and improve robustness. However, given RL methods’ inherent requirement of direct interaction with the environment, and along with No Free Lunch Theorem [9, 10], which suggests a potential trade-off between generalization and specific performance, purely training in the simulation with inaccurate state occupancy distribution will limit the policy’s performance beforehand.
@@
Previous works have attempted to cope with these impediments by reshaping reward functions [11, 12], taking advantage of sample-efficient algorithms [11, 12, 13], or harnessing model-based RL [14]. However, these methods, trained directly in the real world, fail to produce superior locomotion policies and are vulnerable during training compared to zero-shot methods.
@@
Drawing inspiration from how animals use familiar environments to navigate new ones, we investigate to bridge the way back to the controllable simulation environment rather than trapping ourselves in the capricious and noisy real world. This approach offers all-around benefits. In one way, we can sidestep the problem of redesigning a delicate reward function corresponding to the limited observation, and take full advantage of a well-polished simulation training process [1, 3, 4, 15]. Additionally, it requires minimal real-world data while still living off the data abundance available in the simulation.
In this work, we present LoopSR, an effective pipeline for lifelong policy adaptation in legged robots depicted in Fig. 1. At its core, LoopSR utilizes a transformer-based encoder [16] to leverage the latent space. By adopting an autoencoder architecture [17, 18] and contrastive loss [19, 20, 21, 22], the model extracts essential features for reconstructing simulation dynamics, ensuring both effectiveness and generalizability.
We implement learning-based and retrieval-based methods to derive simulation parameters from the latent variable, which compensate for each other in accuracy and robustness to out-of-distribution situations. Through comprehensive evaluations, LoopSR outperforms strong baselines in both the IsaacGym simulated environment [1] and the real-world scene.
@@ II-A Representation Learning in RL
Representation learning [18, 23, 24, 25] has made up the bedrock for modern deep learning algorithms, and the field of RL and control is no exception. Previous works have made versatile attempts to take advantage of the representation learning and latent space [26, 27, 28], including contrastive learning [22, 29], reward function encoding [30, 31], domain/task-specific representation [17, 32, 33], and gait encodings [34, 35], boosting multi-modal input [24] or transferring through different tasks and embodiment [32].
Typically, the domain representation is widely utilized for the task of meta-RL and transfer learning [36, 37]. However, real-world scenarios present a challenge as goal reward functions are not readily available, where we are obliged to focus on learning the representation of the transitions themselves rather than relying on labeled tasks.
@@ II-B Learning in the Real World
Continual learning in RL [38] is a daunting task as the result of the compounding error in drifted distribution. Works have been done to address the problem caused by protean data distribution through experience replay [39, 40], distillation [41], and reusing parts of model [42], etc. However, when it comes to real-world learning, the situation becomes progressively difficult due to noisy input, limited observation, and inaccessible reward functions.
[11, 12, 13] are among the few works tapping into this tricky problem. Instead of sim-to-real transfer, these works have made attempts to completely learn the policy in the real world by introducing reshaped reward functions and techniques of restricting action space to boost effective training. However, these algorithm needs careful designing in action space constraints and rewards to reluctantly obtain effective training and executable gaits, while failing to outstrip zero-shot methods.
With world models exhibiting competence in heterogeneous downstream tasks [43], [14] employ the Dreamer world model in the real-world learning task. The model-based method, though a practical alternative, still suffers from the problem of deficient privileged knowledge. The model’s accuracy in featuring real-world information is also worth questioning in comparison with the simulation, as model-based methods are susceptible to pervasive model biases. The vulnerability of training also places the risk that convergence is not guaranteed facing the composition of model biases and compounding error.
@@ III-B Looping Sim and Real
For pretraining, we employ Proximal Policy Optimization (PPO) [44] in the IsaacGym environment following [3]. The policy is based on DreamWaQ [45] and AMP [46, 47], utilizing an Asymmetric Actor-Critic architecture with expert motion data and domain randomizations [6, 7, 8] to enable real-world deployment. Meanwhile, LoopSR constructs a latent trajectory space by training the trajectory encoder (LoopSR-enc) alongside a multi-head decoder using collected rollouts and privileged information. These rollouts are stored as a reference dataset for later adaptation.
@@
16:     Retrieve $N$ nearest neighbours $z_{1},z_{2},\cdots,z_{N}$ of $z$ in $\mathcal{D}_{z}$, and average the corresponding parameter $c_{1},c_{2},\cdots,c_{N}$ to get $\hat{c}_{1}=[c_{e}^{\prime},c_{r}^{\prime}]$
@@ III-C Network Architecture
To the end of retaining the corresponding parameters of the testing environment, we design the network to cash in the latent space for knowledge distillation from real-world trajectories inspired by the advancements in representation learning methods [17, 18, 48]. The network architecture and its working flow are exhibited in Fig. 2.
Transformer-based Encoder. The key encoder, LoopSR-enc draws on Decision Transformer [49] or specifically follows [50] as the backbone while substituting the embedding of returns-to-go with the one of the next observation. LoopSR-enc $q_{\theta}(z|o_{1:t},a_{1:t},o_{2:t+1})$ receives the total trajectory without reward notation $\mathcal{T}=(o_{1:t},a_{1:t},o_{2:t+1})$ as input. The output latent variable $z\in\mathbb{R}^{32}$ is the average pooling of the latent $z_{i}$ from every timestep $i$ and is modeled as a normal distribution following [18]. To support efficient learning [35, 51], we project $z$ to an $l_{2}$ sphere to bound the latent space.
Autoencoder Architecture. To ensure that $z$ captures enough features from the original trajectory, the training process is modeled as autoencoder(AE)-like with a decoder $q_{\psi}(\hat{o}_{n+1}|z,o_{n},a_{n})$ to reconstruct the new observations $o_{n+1}$ based on the latent variable $z$, the observations $o_{n}$ and the actions $a_{n}$ inspired by [17]. We follow the common approach to use L2 distance as a supplant for the untractable log probability [17, 48, 52] of the next state prediction:
@@
Contrastive Loss. To guide the model to discriminate different terrains implicitly, we apply the contrastive loss in the form of supervised InfoNCE loss similar to [53]. The loss is defined to maximize the inner product of $z$ from the trajectories with the same terrain type label, whose theoretical background of maximizing the mutual information between positive samples and the latent representation is well-discussed in [20, 53].
@@
Figure 2: Illustration of the core network architecture. The whole trajectory is input to a transformer backbone similar to [49], with output $z$ derived from average pooling. The decoders are separately designed to disentangle different traits from the trajectory.
@@ III-E Theoretical Analysis
In (5), $Q^{*}$ and $Q^{*}_{R}$ are optimal value functions in $\mathcal{M}$ and $\mathcal{M}^{R}$ respectively, and $V_{R}^{*}(s^{\prime})$ as the optimal state value function on $\mathcal{M}^{R}$. $\Delta r$ informs the discrepancy in reward functions, which is negligible for the substantiated validity of $r$ from previous works [3, 4, 54, 15]. $\epsilon_{\pi}$ models the discrepancy caused by policy differences, which should approach 00 with sufficient model capacity as $\pi^{*}$ and $\pi^{*}_{R}$ share the same input with detailed reasoning in [55].
@@ IV Experiments
$0.176$ $1.3519\scriptstyle{\pm 0.0003}$ $1.1874\scriptstyle{\pm 0.00043}$ $1.2522\scriptstyle{\pm 0.00023}$ $1.1476\scriptstyle{\pm 0.08996}$ $0.0386\scriptstyle{\pm 0.00506}$ $0.0619\scriptstyle{\pm 0.01521}$
$0.23$ $1.2414\scriptstyle{\pm 0.0354}$ $0.7553\scriptstyle{\pm 0.12505}$ $1.1330\scriptstyle{\pm 0.00813}$ $0.3680\scriptstyle{\pm 0.19475}$ $0.0253\scriptstyle{\pm 0.00024}$ $0.0436\scriptstyle{\pm 0.02800}$
@@
$0.4$ $1.3478\scriptstyle{\pm 0.00012}$ $1.1968\scriptstyle{\pm 0.00087}$ $1.3417\scriptstyle{\pm 0.00020}$ $1.1707\scriptstyle{\pm 0.00694}$ $0.1539\scriptstyle{\pm 0.05367}$ $0.2222\scriptstyle{\pm 0.02665}$
$0.6$ $1.1758\scriptstyle{\pm 0.01412}$ $0.8384\scriptstyle{\pm 0.11217}$ $1.0746\scriptstyle{\pm 0.03184}$ $0.9429\scriptstyle{\pm 0.00661}$ $0.0622\scriptstyle{\pm 0.02096}$ $0.1227\scriptstyle{\pm 0.03877}$
@@
$0.19$ $1.1638\scriptstyle{\pm 0.14744}$ $0.9932\scriptstyle{\pm 0.06514}$ $1.0918\scriptstyle{\pm 0.01214}$ $0.7200\scriptstyle{\pm 0.09696}$ $0.0354\scriptstyle{\pm 0.05214}$ $0.1280\scriptstyle{\pm 0.04706}$
$0.25$ $0.8681\scriptstyle{\pm 0.26708}$ $0.4056\scriptstyle{\pm 0.21618}$ $0.8379\scriptstyle{\pm 0.19660}$ $0.1744\scriptstyle{\pm 0.04831}$ $0.0138\scriptstyle{\pm 0.04024}$ $0.1768\scriptstyle{\pm 0.01260}$
@@ IV-A Sim-to-Sim Experiment
Overview. Although simulation settings cannot perfectly replicate real-world complexities, they offer a controlled environment for conducting statistical experiments, facilitating comparisons when direct real-world computation is impractical. We leverage IsaacGym [1] for training and simulation, and the training code is developed based on [3].
@@
In general, LoopSR runs $100$ sim-to-real loops for each testing environment. In each loop, we collect $31e$ 3 steps in the testing environment and train 500 iterations in the newly constructed training environment. In total, each policy is trained on $51e$ 5 steps of collected data, equivalent to half an hour in a real-world experiment. For a fair comparison, all baselines share the same total training iterations of 80,000. During evaluations, robots are commanded to walk forward at a velocity of $v_{x}^{cmd}=1.0m/s$ to assess their maximum performance. The reward functions follow [3] and are averaged across simulation steps.
@@
  • Expert is a reference policy completely trained on the desired testing environment. The architecture follows DreamWaQ [45].
  • Origin (DreamWaQ) is a reference policy without refinement, representing a zero-shot sim-to-real workflow following DreamWaQ [45].
  • RMA is another zero-shot sim-to-real workflow following RMA [15]. Here, the teacher and student policy are trained for 80,000 iterations respectively.
  • SAC follows [11] with SAC used in pretraining and continual training.
  • World Model indicates a model-based method following [14] with Dreamer [43] as the backbone.
@@ IV-B Sim-to-Real Evaluation
Overview. For real-world evaluations, we utilize the Unitree A1 robot [56] to facilitate real-world deployment. Real-world observation of A1 only includes proprioception $o_{t}\in\mathbb{R}^{45}$ without visual input. The action $a_{t}\in\mathbb{R}^{12}$ specifies target joint positions converted to joint torques via a PD controller. The testing environment consists of diverse real-world terrains, as depicted in Fig. 4.
@@
Slope Up $11.18\scriptstyle{\pm 0.13}$ $11.39\scriptstyle{\pm 0.14}$ $11.59\scriptstyle{\pm 0.09}$

### v3

@@ Abstract
Reinforcement Learning (RL) has shown its remarkable and generalizable capability in legged locomotion through sim-to-real transfer. However, while adaptive methods like domain randomization are expected to enhance policy robustness across diverse environments, they potentially compromise the policy’s performance in any specific environment, leading to suboptimal real-world deployment due to the No Free Lunch theorem. To address this, we propose LoopSR, a lifelong policy adaptation framework that continuously refines RL policies in the post-deployment stage. LoopSR employs a transformer-based encoder to map real-world trajectories into a latent space and reconstruct a digital twin of the real world for further improvement. Autoencoder architecture and contrastive learning methods are adopted to enhance feature extraction of real-world dynamics. Simulation parameters for continual training are derived by combining predicted values from the decoder with retrieved parameters from a pre-collected simulation trajectory dataset. By leveraging simulated continual training, LoopSR achieves superior data efficiency compared with strong baselines, yielding eminent performance with limited data in both sim-to-sim and sim-to-real experiments. Please refer to https://peilinwu.site/looping-sim-and-real.github.io/ for videos and code.
@@ I Introduction
Reinforcement Learning (RL) has seen significant progress as a viable alternative for robotic control, but applying it to real-world scenarios remains challenging. Advances in simulation systems [IsaacGym, mujoco] and extensive parallel training [rudin2022massive] have enhanced policy performance and robustness, but the sim-to-real gap persists.
Adaptive methods such as domain randomization [lee2020challenging, tan2018sim2real, randomization1, randomization2, randomization3] are widely used to address the gap and improve robustness. However, given RL methods’ inherent requirement of direct interaction with the environment, and along with No Free Lunch Theorem [NFL_optim, NFL_supervised], which suggests a potential trade-off between generalization and specific performance, purely training in the simulation with inaccurate state occupancy distribution will limit the policy’s performance beforehand.
@@
Previous works have attempted to cope with these impediments by reshaping reward functions [smith2022walk, smith2022legged], taking advantage of sample-efficient algorithms [smith2022walk, smith2022legged, smith2024grow], or harnessing model-based RL [wu2023daydreamer]. However, these methods, trained directly in the real world, fail to produce superior locomotion policies and are vulnerable during training compared to zero-shot methods.
@@
Drawing inspiration from how animals use familiar environments to navigate new ones, we investigate to bridge the way back to the controllable simulation environment rather than trapping ourselves in the capricious and noisy real world. This approach offers all-around benefits. In one way, we can sidestep the problem of redesigning a delicate reward function corresponding to the limited observation, and take full advantage of a well-polished simulation training process [IsaacGym, rudin2022massive, lee2020challenging, rma]. Additionally, it requires minimal real-world data while still living off the data abundance available in the simulation.
In this work, we present LoopSR, an effective pipeline for lifelong policy adaptation in legged robots depicted in Fig. 1. At its core, LoopSR utilizes a transformer-based encoder [transformer] to leverage the latent space. By adopting an autoencoder architecture [TAE, VAE] and contrastive loss [simsaim, infoNCE, con_gcrl, CURL], the model extracts essential features for reconstructing simulation dynamics, ensuring both effectiveness and generalizability.
We implement learning-based and retrieval-based methods to derive simulation parameters from the latent variable, which compensate for each other in accuracy and robustness to out-of-distribution situations. Through comprehensive evaluations, LoopSR outperforms strong baselines in both the IsaacGym simulated environment [IsaacGym] and the real-world scene.
@@ II-A Representation Learning in RL
Representation learning [VAE, representation_survey, repre_multimodal_sur, repre_con_sur] has made up the bedrock for modern deep learning algorithms, and the field of RL and control is no exception. Previous works have made versatile attempts to take advantage of the representation learning and latent space [repre_rl2, Hilbert, repre_rl], including contrastive learning [CURL, R3M], reward function encoding [reward_encoding, reward_repre_gcrl], domain/task-specific representation [TAE, robot_specific, skill-transfer], and gait encodings [walk_these_ways, Gait_representation], boosting multi-modal input [repre_multimodal_sur] or transferring through different tasks and embodiment [robot_specific].
Typically, the domain representation is widely utilized for the task of meta-RL and transfer learning [transfer_learning2, transfer_learning_survey]. However, real-world scenarios present a challenge as goal reward functions are not readily available, where we are obliged to focus on learning the representation of the transitions themselves rather than relying on labeled tasks.
@@ II-B Learning in the Real World
Continual learning in RL [CRL_definition] is a daunting task as the result of the compounding error in drifted distribution. Works have been done to address the problem caused by protean data distribution through experience replay [CRL_SER, CRL_replay], distillation [CRL_PR_distill], and reusing parts of model [CRL_DET], etc. However, when it comes to real-world learning, the situation becomes progressively difficult due to noisy input, limited observation, and inaccessible reward functions.
[smith2022walk, smith2022legged, smith2024grow] are among the few works tapping into this tricky problem. Instead of sim-to-real transfer, these works have made attempts to completely learn the policy in the real world by introducing reshaped reward functions and techniques of restricting action space to boost effective training. However, these algorithm needs careful designing in action space constraints and rewards to reluctantly obtain effective training and executable gaits, while failing to outstrip zero-shot methods.
With world models exhibiting competence in heterogeneous downstream tasks [dreamer], [wu2023daydreamer] employ the Dreamer world model in the real-world learning task. The model-based method, though a practical alternative, still suffers from the problem of deficient privileged knowledge. The model’s accuracy in featuring real-world information is also worth questioning in comparison with the simulation, as model-based methods are susceptible to pervasive model biases. The vulnerability of training also places the risk that convergence is not guaranteed facing the composition of model biases and compounding error.
@@ III-B Looping Sim and Real
For pretraining, we employ Proximal Policy Optimization (PPO) [PPO] in the IsaacGym environment following [rudin2022massive]. The policy is based on DreamWaQ [dreamwaq] and AMP [AMP1, amp2], utilizing an Asymmetric Actor-Critic architecture with expert motion data and domain randomizations [randomization1, randomization2, randomization3] to enable real-world deployment. Meanwhile, LoopSR constructs a latent trajectory space by training the trajectory encoder (LoopSR-enc) alongside a multi-head decoder using collected rollouts and privileged information. These rollouts are stored as a reference dataset for later adaptation.
@@
16:  Retrieve NN nearest neighbours $z_{1},z_{2},\cdots,z_{N}$ of $z$ in $\mathcal{D}_{z}$, and average the corresponding parameter $c_{1},c_{2},\cdots,c_{N}$ to get $\hat{c}_{1}=[c_{e}^{\prime},c_{r}^{\prime}]$
@@ III-C Network Architecture
To the end of retaining the corresponding parameters of the testing environment, we design the network to cash in the latent space for knowledge distillation from real-world trajectories inspired by the advancements in representation learning methods [TAE, VAE, repre_seq]. The network architecture and its working flow are exhibited in Fig. 2.
Transformer-based Encoder. The key encoder, LoopSR-enc draws on Decision Transformer [DT] or specifically follows [GPT] as the backbone while substituting the embedding of returns-to-go with the one of the next observation. LoopSR-enc $q_{\theta}(z|o_{1:t},a_{1:t},o_{2:t+1})$ receives the total trajectory without reward notation $\mathcal{T}=(o_{1:t},a_{1:t},o_{2:t+1})$ as input. The output latent variable $z\in\mathbb{R}^{32}$ is the average pooling of the latent $z_{i}$ from every timestep $i$ and is modeled as a normal distribution following [VAE]. To support efficient learning [Gait_representation, ASE], we project $z$ to an $l_{2}$ sphere to bound the latent space.
Autoencoder Architecture. To ensure that $z$ captures enough features from the original trajectory, the training process is modeled as autoencoder(AE)-like with a decoder $q_{\psi}(\hat{o}_{n+1}|z,o_{n},a_{n})$ to reconstruct the new observations $o_{n+1}$ based on the latent variable $z$, the observations $o_{n}$ and the actions $a_{n}$ inspired by [TAE]. We follow the common approach to use L2 distance as a supplant for the untractable log probability [TAE, repre_seq, repre_sto_vvp] of the next state prediction:
@@
Contrastive Loss. To guide the model to discriminate different terrains implicitly, we apply the contrastive loss in the form of supervised InfoNCE loss similar to [supervised_con]. The loss is defined to maximize the inner product of $z$ from the trajectories with the same terrain type label, whose theoretical background of maximizing the mutual information between positive samples and the latent representation is well-discussed in [infoNCE, supervised_con].
@@
Figure 2: Illustration of the core network architecture. The whole trajectory is input to a transformer backbone similar to [DT], with output $z$ derived from average pooling. The decoders are separately designed to disentangle different traits from the trajectory.
@@ III-E Theoretical Analysis
In (5), $Q^{*}$ and $Q^{*}_{R}$ are optimal value functions in $\mathcal{M}$ and $\mathcal{M}^{R}$ respectively, and $V_{R}^{*}(s^{\prime})$ as the optimal state value function on $\mathcal{M}^{R}$. $\Delta r$ informs the discrepancy in reward functions, which is negligible for the substantiated validity of $r$ from previous works [rudin2022massive, lee2020challenging, egocentric, rma]. $\epsilon_{\pi}$ models the discrepancy caused by policy differences, which should approach 0 with sufficient model capacity as $\pi^{*}$ and $\pi^{*}_{R}$ share the same input with detailed reasoning in [HIB].
@@ IV Experiments
$0.176$ $3519\scriptstyle{\pm 0.0003} 1.1$$1874\scriptstyle{\pm 0.00043}$ $2522\scriptstyle{\pm 0.00023} 1.1$$1476\scriptstyle{\pm 0.08996}$ $0386\scriptstyle{\pm 0.00506} 0.0$$0619\scriptstyle{\pm 0.01521}$
$0.23$ $1.2414\scriptstyle{\pm 0.0354}$ $0.7553\scriptstyle{\pm 0.12505}$ $1.1330\scriptstyle{\pm 0.00813}$ $3680\scriptstyle{\pm 0.19475} 0.0$$253\scriptstyle{\pm 0.00024} 0.0$$0436\scriptstyle{\pm 0.02800}$
@@
$0.4$ $3478\scriptstyle{\pm 0.00012} 1.1$$1968\scriptstyle{\pm 0.00087}$ $3417\scriptstyle{\pm 0.00020} 1.1$$1707\scriptstyle{\pm 0.00694}$ $0.1539\scriptstyle{\pm 0.05367}$ $0.2222\scriptstyle{\pm 0.02665}$
$0.6$ $1.1758\scriptstyle{\pm 0.01412}$ $0.8384\scriptstyle{\pm 0.11217}$ $1.0746\scriptstyle{\pm 0.03184}$ $9429\scriptstyle{\pm 0.00661} 0.0$$0622\scriptstyle{\pm 0.02096}$ $0.1227\scriptstyle{\pm 0.03877}$
@@
$0.19$ $1.1638\scriptstyle{\pm 0.14744}$ $0.9932\scriptstyle{\pm 0.06514}$ $1.0918\scriptstyle{\pm 0.01214}$ $7200\scriptstyle{\pm 0.09696} 0.0$$0354\scriptstyle{\pm 0.05214}$ $0.1280\scriptstyle{\pm 0.04706}$
$0.25$ $0.8681\scriptstyle{\pm 0.26708}$ $0.4056\scriptstyle{\pm 0.21618}$ $0.8379\scriptstyle{\pm 0.19660}$ $1744\scriptstyle{\pm 0.04831} 0.0$$0138\scriptstyle{\pm 0.04024}$ $0.1768\scriptstyle{\pm 0.01260}$
@@ IV-A Sim-to-Sim Experiment
Overview. Although simulation settings cannot perfectly replicate real-world complexities, they offer a controlled environment for conducting statistical experiments, facilitating comparisons when direct real-world computation is impractical. We leverage IsaacGym [IsaacGym] for training and simulation, and the training code is developed based on [rudin2022massive].
@@
In general, LoopSR runs $100$ sim-to-real loops for each testing environment. In each loop, we collect $1e3$ steps in the testing environment and train 500 iterations in the newly constructed training environment. In total, each policy is trained on $1e5$ steps of collected data, equivalent to half an hour in a real-world experiment. For a fair comparison, all baselines share the same total training iterations of 80,000. During evaluations, robots are commanded to walk forward at a velocity of $v_{x}^{cmd}=1.0m/s$ to assess their maximum performance. The reward functions follow [rudin2022massive] and are averaged across simulation steps.
@@
  • Expert is a reference policy completely trained on the desired testing environment. The architecture follows DreamWaQ [dreamwaq].
  • Origin (DreamWaQ) is a reference policy without refinement, representing a zero-shot sim-to-real workflow following DreamWaQ [dreamwaq].
  • RMA is another zero-shot sim-to-real workflow following RMA [rma]. Here, the teacher and student policy are trained for 80,000 iterations respectively.
  • SAC follows [smith2022walk] with SAC used in pretraining and continual training.
  • World Model indicates a model-based method following [wu2023daydreamer] with Dreamer [dreamer] as the backbone.
@@ IV-B Sim-to-Real Evaluation
Overview. For real-world evaluations, we utilize the Unitree A1 robot [unitree] to facilitate real-world deployment. Real-world observation of A1 only includes proprioception $o_{t}\in\mathbb{R}^{45}$ without visual input. The action $a_{t}\in\mathbb{R}^{12}$ specifies target joint positions converted to joint torques via a PD controller. The testing environment consists of diverse real-world terrains, as depicted in Fig. 4.
@@
Slope Up $1.18\scriptstyle{\pm 0.13} 11$$39\scriptstyle{\pm 0.14} 11.$$1.59\scriptstyle{\pm 0.09}$